# Consensus-based construction of high-dimensional free energy surface

### v3

@@
With the training samples obtained from the aforementioned maximization step, the DNN representation of the FES is optimized using the Adam stochastic gradient descent method [57] for the minimization step. The loss function of the updated DNN representation $A_{\mathcal{N}}(\mathbf{z})$, in turn, navigates the consensus-based adaptive sampling for the updated maximization step. The min-max problem is solved iteratively to achieve comprehensive sampling of the full phase space.
So far, the construction process is based on the function approximation of $A(\mathbf{z})$ over the full regime. However, we note that the kinetic processes of a molecular system are generally characterized by the local minima and saddles points on the FES. On the other hand, the regimes of high free energy are less relevant. To accurately construct these thermodynamically accessible regimes, we modify the loss function $\mathcal{L}$ as
$\mathcal{L}_{\mathcal{N}}(\mathbf{z})=\frac{\left\|\nabla_{\mathbf{z}}A_{\mathcal{N}}(\mathbf{z})+\mathbf{F}(\mathbf{z})\right\|^{2}}{\left\|\mathbf{F}(\mathbf{z})\right\|^{2}+e},$ (14)
for all the biomolecule systems except the toy example of the 1D Rastrigin function. $e$ is a small value to regularize the denominator. We note that a similar formulation is used in Ref. [34] to quantify the model accuracy after the FES is constructed. Alternatively, in this study, $\mathcal{L}_{\mathcal{N}}(\mathbf{z})$ directly involves the construction process for the adaptive sampling of the training set. A detailed algorithm is presented in Algorithm 1.
Algorithm 1 Consensus-based enhance sampling.
Initial sampling point $\mathbf{z}^{i}_{0}$, for $i=1,\ldots,N_{w}$
Initial DNN parameter $\theta_{0}$
The number of training iterations $N_{train}$
The number of data collected $N_{data}$ in each training iteration
@@
         Calculate the mean force $\mathbf{F}^{i}_{t}$ at $\mathbf{z}^{i}_{t}$
@@
         $\mathbf{z}^{i}_{t+1}\leftarrow\mathbf{z}^{i}_{t}-\frac{\alpha}{\gamma}(\mathbf{z}_{t}^{i}-\mathbf{m})\odiv\mathbf{v}+\sqrt{\frac{2\alpha}{\gamma\kappa_{h}}}\mathbf{\eta}_{t},\eta_{t}\sim\mathcal{N}(0,1)$
@@
     Save the training dataset $\mathcal{D}_{j}=\{\mathbf{z}^{i}_{t},\mathbf{F}^{i}_{t}\}_{t=0}^{T}$
@@
Numerical examples
One-dimensional Rastrigin function
@@
To illustrate the essential idea of the present method, we start with the 1-dimensional Rastrigin function:
$f(x)=x^{2}-\cos(2\pi x),\quad x\in[-3,3].$ (15)
@@
Figure 2: Adaptive sampling and construction of the 1D Rastrigin function. Left: The reference function $f(x)$ and the constructed approximations $f_{\theta_{i}}(x)$ obtained at different iteration steps. The relative $l_{2}$ error is less than $6\times 10^{-3}$ after $12$ iterations. Right: The residual function $|f(x)-f_{\theta_{i}}(x)|$. The symbols represent the locations identified by each adaptive sampling (i.e., the maximization) step where new points will be added for the next training (i.e., the minimization) step.
Instead of a neural network, we simply use a piecewise polynomial function $f_{\theta}(x)$ for this 1D problem, where $\theta$ represents the fitting parameters. Accordingly, the loss function is directly defined as $\|f-f_{\theta_{i}}\|$. Initially, we set $f_{\theta_{0}}(x)\equiv 8$ with consistent boundary condition. We use the proposed sampling method with 10 walkers to estimate the first and second moments, yielding $m_{1}=1\times 10^{-4}$ and $V_{1}=0.022$ and therefore the first and second derivative $f^{\prime}(m_{1})=0$ and $f^{\prime\prime}(m_{1})=\frac{1}{V_{1}}=40.18$. This value is extremely close to the actual minimizer at 00 (i.e., the max-residue point) and the second derivative $f^{\prime\prime}(m_{1})=41.08$. Accordingly, we add new data points near $x=1\times 10^{-4}$ into the training set consisting of the boundary points $f(-3)=f(3)=8$, which enables us to construct an updated approximation $f_{\theta_{1}}(x)$. Similarly, with the approximation $f_{\theta_{i-1}}(x)$, we conduct another sampling process (i.e., the maximization step) and include the obtained samples near $m_{i}$ into the training set, yielding an updated approximation $f_{\theta_{i}}(x)$. As shown in Figure 2, for each iteration, the sampling step can pinpoint the regime where the error is most pronounced. The underlying function $f(x)$ can be fully recovered after $12$ iterations.
Two-dimensional FES
Figure 3: The 2D FES for the Ala2 molecule on the $\phi-\psi$ plane constructed by (a) Metadynamics (reference) (b) VES (c) RiD (d) CES. The accuracy and the computational cost are shown in Table 1.
@@
We use the alanine dipeptide (Ace-Ala-Nme), referred to as Ala2, as a benchmark problem. The molecule is solvated in $383$ TIP3P water molecules similar to Ref. [42]. The full MD system is simulated in a canonical ensemble under temperature $300K$ using the Amber99-SB force field [58] with a time step is 2 fs. We refer to the Appendix for the simulation details.
We choose the CVs as the torsion angles: $\phi$ (C, N, Cα, C) and $\psi$ (N, Cα, C, N). For comparison, we also construct the FESs using the VES  [8] and RiD  [42] method with the same setup parameters presented therein. For the CES method, we use $10$ walkers to explore the configuration space. The initial points of the walkers at the $k$th iteration are chosen to be the final $10$ points of a biased dynamics with the bias potential $-A_{\mathcal{N}}(\mathbf{z};\theta_{k-1})$. In the sampling stage, the inverse of the low and high temperatures are set to be $\kappa_{l}=10$ and $\kappa_{h}=1$, respectively. For each sample point, restrained dynamics is conducted with $\kappa=500$ for $5000$ steps to compute the average force. The timestep $\alpha$ for the dynamics of the random walkers (11) is set to be $0.1$. Figure 3 shows the FESs constructed by the three different methods and the reference obtained by the metadynamics [4] using a long simulation time. As presented in Table 1, the CES method yields smaller approximation error and meanwhile requires lower computational cost. The better performance is not unexpected since the sample points can be adaptively optimized based on the construction error.
It is worth mentioning that the VES method was initially developed to achieve efficient enhanced sampling in the high-dimensional phase space; it is somewhat unfair to use it for direct FES construction since the estimation of the probability density function from the obtained samples could be numerically challenging for high-dimensional cases. On the other hand, the present CES method enables the explicit construction of FES if the analytical form is needed. Also, we note that the accuracy of the RiD method can be further improved using more iterations at the cost of larger simulation and training overhead. Below, we test the methods in more complex systems.
@@
Table 1: The accuracy of the constructed 2D FES (the Ala2 molecule) and computational time (in hours, the same below) for the VES, RiD, and CES methods. The $l_{2}$ and $l_{\infty}$ error are computed up to $40$ KJ/mol. The reference solution is constructed by the metadynamics. The simulation time of the CES method is multiplied by $10$ since 10 walkers are used.
Three-dimensional FES
Next, we consider a s-(1)-phenylethyl (s1pe) peptoid in an aqueous environment similar to Refs. [59, 53]. The full system consists of one biomolecule and 546 water molecules in a $(2.9{\rm nm})^{3}$ dodecahedron box. The CHARMM general force field (CGenFF) [59] is used for the biomolecule and the TIP3P model [60] is used for the water molecules. The system temperature is set to be 298 K with a time step is $2$ fs. We refer to the Appendix for details.
The CVs are the three torsion angles $\omega$ (Cα, C, N, Cα), $\phi$, and $\psi$, where the latter two are the same as the Ala2 molecule. The FES is constructed by both the RiD and CES methods. The setup of the RiD method is the same as [53]. For the CES method, we use 20 walkers and set the inverse of the low and high temperatures to be $\kappa_{l}=10$ and $\kappa_{h}=2$, respectively. The initial points of these walkers at each iteration are chosen in the same method as the two-dimensional problem. The timestep $\alpha$ for the dynamics of the random walkers (11) is set to be $0.1$. For each sample point, restrained dynamics with $\kappa=500$ is conducted for $10000$ steps to compute the mean force.
For visualization, we project the constructed FES onto a two-dimensional plane and fix the third variable. Figure 4 shows the projected FES on the $\omega-\phi$ and $\phi-\psi$ plane obtained from the CES and the RiD methods. For each projection, the reference is constructed as a 2D FES using the metadynamics [4]. Similar to the previous 2D case, the present CES method yields higher accuracy with lower computational cost.
Figure 4: The 3D FES for molecule s1pe projected on the $\omega-\phi$ with $\psi=1.5$ (left column) and $\phi-\psi$ planes with $\omega=1.5$ (right-column). (a-b) The 2D FES constructed using metadynamics with the third variable restrained (reference); (c-d) projection of the 3D FES constructed by the RiD method; (e-f) projection of the 3D FES constructed by the CES method.
Method Restraint Accuracy Time
$l_{2}$ error $l_{\infty}$ error Simulation Train
RiD $\psi=1.5$ $5.76$ $25.72$ $423.33$ $8$ (GPU)
$\omega=1.5$ $12.04$ $49.13$
CES $\psi=1.5$ $2.44$ $11.21$ $4.81\times 20$ $0.84$ (GPU)
$\omega=1.5$ $3.89$ $28.80$
Table 2: The accuracy of the constructed 3D FES (the s1pe molecule) and computational time for RiD and CES methods. The $l_{2}$ and $l_{\infty}$ error are calculated up to $40$ KJ/mol. For each case, the FES is projected onto a 2D plane with the third variable fixed; the reference solution is constructed as a 2D FES by the metadynamics. The simulation time of the CES method is multiplied by $20$ since 20 walkers are used.
Nine-dimensional FES
Furthermore, we consider a more complex molecule, the peptoid trimer (s1pe)3, solvated in a (4.2 nm)3 dodecahedron box with 1622 TIP3P water molecules. The force field and other simulation setups are similar to the s1pe molecule.
The chosen CVs consist of the $9$ torsion angles $\omega,\phi,\psi$ that are defined in Ala2 and s1pe case, associated with the different Cα atoms and denoted as $\omega_{1},\phi_{1},\psi_{1},\omega_{2},\phi_{2},\psi_{2},\omega_{3},\phi_{3},\psi_{3}$. We use 64 walkers for this case and the initial conditions of the walkers at each iteration are chosen with the same method as the previous cases. The inverse of the low and the high temperature are set to be $\kappa_{l}=100$ and $\kappa_{h}=2$, respectively. The timestep $\alpha$ of the dynamics of the random walkers is set to be $0.1$. The FES is constructed by the CSE method using 28 iterations of sampling and training, which requires $225.53\times 64=14434.35$ CPU hours for simulation and $6.06$ GPU hours for training. For comparison, the RiD method uses $17900$ CPU hours for simulation and $15.44$ GPU hours for training. Similar to the above 3D problem, the constructed 9D FES is projected onto various 2D planes with the remaining variables fixed. For each 2D projection, the reference is constructed as a 2D FES using metadynamics. Fig. 5 shows the projection on the $\omega_{1}-\phi_{1}$ and $\omega_{1}-\psi_{1}$ plane (see Appendix for visualization of other 2D projections). Compared with the Rid method, the present CES method yields higher accuracy with lower computational cost.
The different performances could be possibly due to the distinct ways of imposing the sampling adaptivity for the two methods. Specifically, the Rid method trains a replica of DNNs on the same sample set and uses the standard deviation of multiple DNNs’ predictions as an indirect measure of the construction error. The significant sampling error of the mean force (i.e., $-\nabla A(\mathbf{z})$) could be overlooked by using the standard deviation as the uncertainty indicator, and further propagate into the construction error. For instance, a biased sampling of mean force may lead to a consistent biased prediction among multiple DNNs. As a result, even if the standard deviation of multiple DNNs’ predictions for a sample point is small, the construction error could be still pronounced in that region. Instead, the CES method directly uses the construction error to impose the adaptivity and therefore achieve enhanced sampling of the regions lacking accuracy.
Figure 5: The 9D FES for molecule (s1pe)3 projected on the $\omega_{1}-\phi_{1}$ (left column) and $\omega_{1}-\psi_{1}$ plane (right column). (a-b) The 2D FES constructed using metadynamics with the remaining variable restrained (reference); (c-d) projection of the 9D FES constructed by the RiD method using $17900$ CPU hours for simulation and $15.44$ GPU hours for training; (e-f) projection of the 9D FES constructed by the CES method using $14434.35$ CPU hours for simulation and $6.06$ GPU hours for training.
Thirty dimensional FES
Finally, we consider polyalanine-15 (Ace-(Ala)15-Nme), denoted as Ala16, solvated in $2258$ water molecules. The full system is simulated temperature 300K in a ($4.62$ nm)3 dodecahedron box using the Amber99-SB force field with a time step is 2 fs. We refer to the Appendix for the simulation details.
The chosen CVs consist of torsion angles $\phi$ and $\psi$, defined in Ala2 case, associated with the different Cαs, denoted as $\{\phi_{i},\psi_{i}\}_{i=1}^{15}$. We use 64 walkers for this case and the initial value of the walkers at each iteration is given by the biased simulation before. The inverse of the low and high temperatures are set to be $\kappa_{l}=20$ and $\kappa_{h}=5$, respectively. The timestep $\alpha$ for dynamics of the random walkers 11 is set to be $0.1$. The FES is constructed using $100$ iterations of sampling and training. Similar to the previous case, the obtained FES is plotted on a two-dimensional plane while the remaining variables are fixed and the reference is constructed as a 2D FES using the metadynamics. Fig. 6 shows the projection on the $\phi_{1}-\psi_{1}$, $\phi_{2}-\phi_{3}$ and $\psi_{2}-\psi_{3}$ plane. For all the cases, the projection of the 30-dimensional FES shows good agreement with the 2-dimensional reference solution. We have also examined the projection on other planes; the prediction shows good agreement with the reference solution as well. We refer to the Appendix for details.
@@
Figure 6: The 30D FES for molecule Ala16 projected on various 2D planes. (a-b) $\phi_{1}-\psi_{1}$ (c-d) $\phi_{2}-\phi_{3}$ (e-f) $\psi_{2}-\psi_{3}$. (a-c-e) 2D FES constructed by metadynamics (reference) (b-d-f) The 30D FES constructed by the present CES method.
Conclusion

### v4

@@ 2.5 Practical stabilization modifications
Furthermore, we note that the kinetic processes of a molecular system are generally characterized by the local minima and saddle points of the FES. On the other hand, the regimes of high free energy are less relevant. To accurately construct these thermodynamically accessible regimes, we modify the loss function as
(28) $\mathcal{L}_{\mathcal{N}}(\mathbf{z})=\frac{\left|\nabla_{\mathbf{z}}A_{\mathcal{N}}(\mathbf{z})+\mathrm{F}(\mathbf{z})\right|^{2}}{\left|\mathrm{F}(\mathbf{z})\right|^{2}+e},$
for all the biomolecule systems except the toy example of the 1D Rastrigin function. $e$ is a small value to regularize the denominator.
With the training samples obtained from the aforementioned maximization step, the NN surrogate is optimized using the Adam stochastic gradient descent method [25] for the minimization step. The loss function of the updated $A_{\mathcal{N}}(\mathbf{z})$, in turn, navigates the consensus-based adaptive sampling for the updated maximization step. The min-max problem is solved iteratively to achieve comprehensive sampling of the full phase space. A detailed algorithm is presented in Alg. 1.
Algorithm 1 Consensus-based adaptive sampling.
0:  Initial sampling point $\mathbf{z}^{i}_{0}$, for $i=1,\ldots,N_{w}$
0:  Initial NN parameter $\theta_{0}$
0:  The number of training iterations $N_{train}$
0:  The number of data collected $N_{data}$ in each training iteration
@@
        calculate the mean force $\mathrm{F}^{i}_{t}$ at $\mathbf{z}^{i}_{t}$
@@
        $\mathbf{z}^{i}_{t+1}\leftarrow\mathbf{z}^{i}_{t}-\frac{\delta t}{\gamma}(\mathbf{z}_{t}^{i}-\mathbf{m})\odiv\mathbf{v}+\sqrt{\frac{2\delta t}{\gamma\kappa_{h}}}\mathbf{\eta}^{i}_{t},\eta^{i}_{t}\sim\mathcal{N}(0,1)$
@@
     Save the training dataset $\mathcal{D}_{j}=\{\mathbf{z}^{i}_{t},\mathrm{F}^{i}_{t}\}_{t=0}^{T}$
@@
2.6 Related methods
Before we present the numerical results, we further discuss the differences between the present CAS method and the existing approaches. Since the standard sampling methods such as MCMC [37], Langevin dynamics [38], and generative model [12, 42] are not applicable, we focus on two major approaches based on the enhanced sampling such as the VES method [4] as well as the active learning such as the RiD method [54].
The VES method provides an efficient approach to establish the re-weighted sampling of the phase space. However, the surrogate construction of the FES $A(\mathbf{z})$ further relies on the estimation of the PDF $\rho(\mathbf{z})$ from the obtained reweighted samples, which could become computationally expensive for high-dimensional problems. In contrast, the present method circumvents the PDF estimation and directly constructs $A(\mathbf{z})$ through the adaptive sampling of the mean force $F(\mathbf{z})=-\nabla A(\mathbf{z})$.
Furthermore, the present method differs from the RiD method in the following two aspects: (\@slowromancapi@) The present method imposes the sampling adaptivity by explicitly using the residual error $\mathcal{L_{N}}(\mathbf{z})$ while the RiD method does not directly account for the approximation error. Instead, the RiD method trains a replica of NNs on the same sample set and uses the standard deviation of multiple NN predictions as an indirect measure of the $\mathcal{L_{N}}(\mathbf{z})$. As a result, the sampling error of the mean force $F(\mathbf{z})$ may lead to a consistent biased prediction among multiple NNs with a small standard deviation. (\@slowromancapii@) The RiD method drives the sampling dynamics with a biased MD potential $U(\mathbf{r})$ similar to Eq. (3). Accordingly, the time step needs to be sufficiently small due to the stiffness of $U(\mathbf{r})$. In contrast, the dynamics of the random walkers for the present method is governed by a mean-field quadratic potential $G(\mathbf{z})$ in Eq. (12). In particular, $G(\mathbf{z})$ is decoupled from the MD potential $U(\mathbf{r})$. This enables us to choose a much larger time step irrespective of the roughness of $U(\mathbf{r})$ and achieve a more efficient exploration of the phase space. Below we compare the performance of the present method with these two methods in more detail.
3 Numerical Results
3.1 One-dimensional Rastrigin function
@@
To illustrate the essential idea of the present method, we start with the one-dimensional Rastrigin function:
(29) $f(z)=z^{2}-\cos(2\pi z),\quad z\in[-3,3].$
[…]
@@
Figure 2: Adaptive sampling and construction of the 1D Rastrigin function. Left: The reference function $f(z)$ and the constructed approximations $f_{\theta_{i}}(z)$ obtained at different iteration steps. The relative $l_{2}$ error is less than $6\times 10^{-3}$ after $12$ iterations. Right: The residual function $|f(z)-f_{\theta_{i}}(z)|$. The symbols represent the locations identified by each adaptive sampling (i.e., the maximization) step where new points will be added for the next training (i.e., the minimization) step.
3.2 Two-dimensional FES
@@
We use the alanine dipeptide (Ace-Ala-Nme), referred to as Ala2, as a benchmark problem. The molecule is solvated in $383$ TIP3P water molecules similar to Ref. [54]. The full MD system is simulated in a canonical ensemble under temperature $300$K using the Amber99-SB force field [22] with a time step $2$ fs. We refer to Appendix A for the simulation details.
[…]
Table 1: The accuracy of the constructed 2D FES (the Ala2 molecule) and computational time (in hours, the same below) for the VES, RiD, and CAS methods. The $l_{2}$ and $l_{\infty}$ error are computed up to $40$ KJ/mol. The reference solution is constructed by the metadynamics. The simulation time of the CAS method is multiplied by $10$ since 10 random walkers are used.
@@
Figure 3: The 2D FES for the Ala2 molecule on the $\phi-\psi$ plane constructed by (a) Metadynamics (reference) (b) VES (c) RiD (d) CAS (the present method). The accuracy and the computational cost are shown in Table 1.
3.3 Three-dimensional FES
Next, we consider a s-(1)-phenylethyl (s1pe) peptoid in an aqueous environment similar to Refs. [50, 48]. The full system consists of one biomolecule and 546 water molecules in a $(2.9{\rm nm})^{3}$ dodecahedron box. The CHARMM general force field (CGenFF) [50] is used for the biomolecule and the TIP3P model [24] is used for the water molecules; see Appendix A for details.
The CVs are the three torsion angles $\omega$ spanned by atoms (Cα, C, N, Cα), $\phi$, and $\psi$, where the latter two are the same as the Ala2 molecule. The FES is constructed by both the RiD and CAS methods. The setup of the RiD method is the same as [48]. For the CAS method, we use 20 walkers and set the inverse of the low and high temperatures to be $\kappa_{l}=10$ and $\kappa_{h}=2$, respectively. The initial points of these walkers at each iteration are chosen in the same method as the two-dimensional problem. The timestep $\delta t$ for the dynamics of the random walkers (12) is set to be $0.1$. For each sample point, restrained MD with $\kappa=500$ is conducted for $10000$ steps to compute the mean force.
For visualization, we project the constructed FES onto a two-dimensional plane and fix the third variable. Fig. 4 shows the projected FES on the $\omega-\phi$ and $\phi-\psi$ plane obtained from the CAS and the RiD methods. For each projection, the reference is constructed as a 2D FES using the metadynamics [27]. Similar to the previous 2D case, the present CAS method yields higher accuracy with lower computational cost.
Figure 4: The 3D FES for molecule s1pe projected on the $\omega-\phi$ plane by fixing $\psi=1.5$ (first row) and the $\phi-\psi$ planes by fixing $\omega=1.5$ (second-row). (a-d) The 2D FES constructed using metadynamics with the third variable restrained (reference); (b-e) 2D projection of the 3D FES constructed by the RiD method; (c-f) 2D projection of the 3D FES constructed by the present CAS method.
3.4 Nine-dimensional FES
Figure 5: The 9D FES for molecule (s1pe)3 projected on the $\omega_{1}-\phi_{1}$ (first row) and $\omega_{1}-\psi_{1}$ plane (second row). (a-d) The 2D FES constructed using metadynamics with the remaining variable restrained (reference solution); (b-e) 2D projection of the 9D FES constructed by the RiD method; (c-f) 2D projection of the 9D FES constructed by the present CAS method.
Furthermore, we consider a more complex molecule, the peptoid trimer (s1pe)3, solvated in a (4.2 nm)3 dodecahedron box with 1622 TIP3P water molecules. The chosen CVs consist of the $9$ torsion angles $\omega,\phi,\psi$ that are defined in s1pe case, associated with the different Cα atoms and denoted as $\omega_{1},\phi_{1},\psi_{1},\omega_{2},\phi_{2},\psi_{2},\omega_{3},\phi_{3},\psi_{3}$. We use 64 walkers for this case and the initial conditions of the walkers at each iteration are chosen with the same method as the previous cases. The inverse of the low and the high temperature are set to be $\kappa_{l}=100$ and $\kappa_{h}=2$, respectively. The timestep $\delta t$ of the dynamics of the random walkers is set to be $0.1$. The FES is constructed by the CAS method using 28 iterations of sampling and training, which requires $225.53\times 64=14434.35$ CPU hours for simulation and $6.06$ GPU hours for training. For comparison, the RiD method uses $17900$ CPU hours for simulation and $15.44$ GPU hours for training. Similar to the above 3D problem, the constructed 9D FES is projected onto various 2D planes with the remaining variables fixed. For each 2D projection, the reference is constructed as a 2D FES using metadynamics. Fig. 5 shows the projection on the $\omega_{1}-\phi_{1}$ and $\omega_{1}-\psi_{1}$ plane (see Appendix C.4 for other 2D projections). Compared with the Rid method, the present CAS method yields higher accuracy and lower computational cost.
3.5 Thirty-dimensional FES
Finally, we consider the polyalanine-15 (Ace-(Ala)15-Nme), denoted as Ala16, solvated in $2258$ water molecules. The chosen CVs consist of torsion angles $\phi$ and $\psi$, defined in Ala2 case, associated with the different Cαs, denoted as $\{\phi_{i},\psi_{i}\}_{i=1}^{15}$. We use 64 walkers for this case and the initial value of the walkers at each iteration is given by the biased simulation before. The inverse of the low and high temperatures are set to be $\kappa_{l}=20$ and $\kappa_{h}=5$, respectively. The timestep $\alpha$ for dynamics of the random walkers Eq. (12) is set to be $0.1$. The FES is constructed using $100$ iterations of sampling and training. Similar to the previous case, the obtained FES is plotted on a two-dimensional plane while the remaining variables are fixed and the reference is constructed as a 2D FES using the metadynamics. Fig. 6 shows the projection on the $\phi_{1}-\psi_{1}$, $\phi_{2}-\phi_{3}$ and $\psi_{2}-\psi_{3}$ plane. For all the cases, the projection of the 30-dimensional FES shows good agreement with the 2-dimensional reference solution. We have also examined the projection on other planes; the prediction shows good agreement with the reference solution as well.
Figure 6: The 30D FES for molecule Ala16 projected on various 2D planes. (a-d) $\phi_{1}-\psi_{1}$ (b-e) $\phi_{2}-\phi_{3}$ (c-f) $\psi_{2}-\psi_{3}$. (a-b-c) 2D FES constructed by metadynamics (reference); (d-e-f) 2D projection of the 30D FES constructed by the present CAS method.
4 Conclusions